Use the Laplace transform to solve the given initial-value problem.




y′ + 2y = sin 4t, y(0) = 1

1
Expert's answer
2022-05-23T23:10:29-0400

Solution

Let L(y) = Y(s)

Then L(y’) = sY(s) – y(0) = sY – 1; L(sin(4t) = 4/(s2 + 42)

So from the given initial-value problem

sY + 2Y – 1 = 4/(s2 + 42)

Y = 4/[(s2 + 42)(s + 2)] + 1/(s + 2)

4/[(s2 + 42)(s + 2)] = (As+B)/(s2 + 42) + C/(s + 2) => (As+B)(s+2) + C(s2 + 42) = 4 => As2 +(2A+B)s + 2B +Cs2 +16C = 4 => A + C = 0, 2A + B = 0, 2B + 16C = 4 => A = -C, B – 2C = 0, B + 8C = 2 => C = 0.2, B = 0.4, A = -0.2 =>

Y(s) = 0.2(2 – s)/(s2 + 42) + 1.2/(s + 2) = 0.4/(s2 + 42) – 0.2s/(s2 + 42) + 1.2/(s + 2)

Therefore using tables of inverse Laplace transform

y(t) = L-1[Y(s)] = 0.1sin(4t) – 0.2 cos(4t) + 1.2e-2t  

Answer

y(t) = 0.1sin(4t) – 0.2 cos(4t) + 1.2e-2t